OpenStudy (anonymous):

Lucas cut a board that is 3 1/5 ft long and 3 1/3 ft wide. What is the area of the board? Express in simplest form

OpenStudy (tgannon998):

What do you think it is?

OpenStudy (anonymous):

A. 6 8/15 B. 9 1/15 C. 9 2/3 D. 10 2/3

OpenStudy (anonymous):

@flower2468

OpenStudy (tgannon998):

D

OpenStudy (anonymous):

No direct answers!

OpenStudy (anonymous):

and I need someone to explain it to me! thxs

OpenStudy (tgannon998):

Multiply the sides together

OpenStudy (anonymous):

@rebeccaxhawaii

OpenStudy (anonymous):

@Qwertty123

OpenStudy (anonymous):

@phi

OpenStudy (phi):

can you change 3 1/5 to an improper fraction ?

OpenStudy (anonymous):

no

OpenStudy (phi):

first change 3 by multiplying it by 5/5 what do you get ?

OpenStudy (mathmale):

Area of a rectangle = A = (length) * (width). Here, the length is ... ? and the width is ... ? So the area is (length) * (width) = ?

OpenStudy (anonymous):

25

OpenStudy (mathmale):

And yes, the first step is to convert these mixed numbers to improper fractions.

OpenStudy (phi):

what is 3*5 / 5 ?

OpenStudy (anonymous):

25

OpenStudy (phi):

what is 3*5 ? ( not 25!)

OpenStudy (anonymous):

15 i meant

OpenStudy (phi):

what is 3*5 / 5 ?

OpenStudy (anonymous):

sorry i got mixed up with 5 times 5

OpenStudy (anonymous):

3

OpenStudy (phi):

yes, 3 is the same as 15/5 (if you simplify it you get back 3) anyway 3 1/5 is the same as \[ \frac{15}{5}+\frac{1}{5} \] you add the tops and keep the same bottom

OpenStudy (anonymous):

3 1/5

OpenStudy (phi):

ok, but we are changing 3 1/5 to 15/5 + 1/5 to get an improper fraction can you add \[ \frac{15}{5} + \frac{1}{5}\]?

OpenStudy (anonymous):

3 1/5

OpenStudy (phi):

you add fractions with the *same bottom*, you add their tops

OpenStudy (anonymous):

???

OpenStudy (phi):

what is 15+1 ?

OpenStudy (anonymous):

16

OpenStudy (phi):

that means \[ \frac{15}{5}+\frac{1}{5}= \frac{16}{5} \] what you did is change 3 1/5 to 16/5

OpenStudy (anonymous):

yes

OpenStudy (phi):

people learn a "rule"